OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 110r3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

ganeshie8 (ganeshie8):

Each piece has an area of square centimeter somehting is missing at above line ?

ganeshie8 (ganeshie8):

also, area is it \[ 110r^3 \] ?

OpenStudy (anonymous):

yes an the part that is missing is in the picture

OpenStudy (anonymous):

????

ganeshie8 (ganeshie8):

\[ \frac{100r^3}{ \sqrt{r^5}}\]

ganeshie8 (ganeshie8):

thats the required expression. its just ration of "total area of card board" and "area of each paper piece" you can simplify it further by pulling the r up.. have fun

OpenStudy (anonymous):

umm? ok

OpenStudy (anonymous):

You have an area of size 110r^3 square centimeters. You have Y number of pieces that cover that area. Each piece is (sqrt(r^5)) square centimeters. Basically you're stating that Y number of pieces of a particular size = the total area So: \[Y \sqrt{r ^{5}} = 110r ^{3}\] \[Y = \frac{110r ^{3}}{\sqrt{r ^{5}}}\] Simplify for R.